Expand ....
(x-3)^2 - (x+3)^2

and Expand
(x-2)(3x^2+2x-1)

Answer by ewatrrr(24785)

 
Hi
(x-3)^2 - (x+3)^2 = [x^2-6x+9] - [x^2+6x-9] = -12x + 18
(x-2)(3x^2+2x-1) = x(3x^2+2x-1)-2(3x^2+2x-1)= 3x^3-4x^2-5x + 2